At a hardware store buying 20 fence pickets and 4 posts cost 92 dollars,while 30 fence pickets and 5 posts cost 130 dollars.Find the price of each fence picket and the price of each post.
Answer by checkley79(3341)

20F+4P=92 MULTIPLY BY -1.5 & ADD.
30F+5P=130
-30F-6P=-138
-------------
-P=-8
P=$8 COST OF ONE POST.
20F+4*8=92
20F+32=92
20P=92-32
20P=60
P=$3 COST PER FENCE PICKET.
PROOF:
30*3+5*8=130
90+40=130
130=130